# A Comprehensive Analysis on LLM-based Node Classification Algorithms

### v1

@@ Abstract
Node classification is a fundamental task in graph analysis, with broad applications across various fields. Recent breakthroughs in Large Language Models (LLMs) have enabled LLM-based approaches for this task. Although many studies demonstrate the impressive performance of LLM-based methods, the lack of clear design guidelines may hinder their practical application. In this work, we aim to establish such guidelines through a fair and systematic comparison of these algorithms. As a first step, we developed LLMNodeBed, a comprehensive codebase and testbed for node classification using LLMs. It includes ten datasets, eight LLM-based algorithms, and three learning paradigms, and is designed for easy extension with new methods and datasets. Subsequently, we conducted extensive experiments, training and evaluating over 2,200 models, to determine the key settings (e.g., learning paradigms and homophily) and components (e.g., model size) that affect performance. Our findings uncover eight insights, e.g., (1) LLM-based methods can significantly outperform traditional methods in a semi-supervised setting, while the advantage is marginal in a supervised setting; (2) Graph Foundation Models can beat open-source LLMs but still fall short of strong LLMs like GPT-4o in a zero-shot setting. We hope that the release of LLMNodeBed, along with our insights, will facilitate reproducible research and inspire future studies in this field. Codes and datasets are released at https://llmnodebed.github.io/.
@@ 1 Introduction
Given its practical importance, node classification has been a long-standing research focus in both academia and industry. The earliest attempts to address this task adopted techniques such as Laplacian regularization (Belkin et al., 2006), graph embeddings (Yang et al., 2016a), and label propagation (Zhu et al., 2003). Over the past decade, GNN-based methods have been developed and have quickly become prominent due to their superior performance, as demonstrated by works such as Kipf & Welling (2017), Veličković et al. (2018), and Hamilton et al. (2017). Additionally, the incorporation of encoded textual information has been shown to further complement GNNs’ node features, enhancing their effectiveness (Jin et al., 2023b; Zhao et al., 2023a).
@@
Inspired by the recent success of LLMs, there has been a surge of interest in leveraging LLMs for node classification (Li et al., 2023). LLMs, pre-trained on extensive text corpora, possess context-aware knowledge and superior semantic comprehension, overcoming the limitations of the non-contextualized shallow embeddings used by traditional GNNs. Typically, supervised methods fall into three categories: Encoder, Reasoner, and Predictor. In the Encoder paradigm, LLMs employ their vast parameters to encode nodes’ textual information, producing more expressive features that surpass shallow embeddings (Zhu et al., 2024b). The Reasoner approach utilizes LLMs’ reasoning capabilities to enhance node attributes and the task descriptions with a more detailed text (Chen et al., 2024b; He et al., 2023). This generated text augments the nodes’ original information, thereby enriching their attributes. Lastly, the Predictor role involves LLMs integrating graph context through graph encoders, enabling direct text-based predictions (Chen et al., 2024a; Tang et al., 2023; Chai et al., 2023; Huang et al., 2024). For zero-shot learning with LLMs, methods can be categorized into two types: Direct Inference and Graph Foundation Models (GFMs). Direct Inference involves guiding LLMs to directly perform classification tasks via crafted prompts (Huang et al., 2023). In contrast, GFMs entail pre-training on extensive graph corpora before applying the model to target graphs, thereby equipping the model with specialized graph intelligence (Li et al., 2024b). An illustration of these methods is shown in Figure 2.
@@
Despite tremendous efforts and promising results, the design principles for LLM-based node classification algorithms remain elusive. Given the significant training and inference costs associated with LLMs, practitioners may opt to deploy these algorithms only when they provide substantial performance enhancements compared to costs. This study, therefore, seeks to identify (1) the most suitable settings for each algorithm category, and (2) the scenarios where LLMs surpass traditional LMs such as BERT. While recent work like GLBench (Li et al., 2024c) has evaluated various methods using consistent data splits in semi-supervised and zero-shot settings, differences in backbone architectures and implementation codebases still hinder fair comparisons and rigorous conclusions. To address these limitations, we introduce a new benchmark that further standardizes backbones and codebases. Additionally, we extend GLBench by incorporating three new E-Commerce datasets relevant to practical applications and expanding the evaluation settings. Specifically, we assess the impact of supervision signals (e.g., supervised, semi-supervised), different language model backbones (e.g., RoBERTa, Mistral, LLaMA, GPT-4o), and various prompt types (e.g., CoT, ToT, ReAct). These enhancements enable a more detailed and reliable analysis of LLM-based node classification methods. In summary, our contributions to the field of LLMs for graph analysis are as follows:
@@
  • A Testbed: We release LLMNodeBed, a PyG-based testbed designed to facilitate reproducible and rigorous research in LLM-based node classification algorithms. The initial release includes ten datasets, eight LLM-based algorithms, and three learning configurations. LLMNodeBed allows for easy addition of new algorithms or datasets, and a single command to run all experiments, and to automatically generate all tables included in this work.
@@
  • Comprehensive Experiments: By training and evaluating over 2,200 models, we analyzed how the learning paradigm, homophily, language model type and size, and prompt design impact the performance of each algorithm category.
@@ 2 Preliminaries on Node Classification
To leverage the language abilities of LLMs, we study the node classification task within the context of text-attributed graphs (TAGs) (Ma & Tang, 2021). TAGs are represented as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{S})$, where $\mathcal{V}$ represents the set of nodes, $\mathcal{E}$ the set of edges, and $\mathcal{S}$ the collection of textual descriptions associated with each node $v\in\mathcal{V}$. Some of the nodes are associated with labels, represented as $\mathcal{V}_{l}\subset\mathcal{V}$. The remaining nodes do not have labels, and are denoted as $\mathcal{V}_{u}$. The goal of node classification is to train a neural network based on the graph $\mathcal{G}$ and the labels of $\mathcal{V}_{l}$, which can predict the labels of the unlabeled nodes in $\mathcal{V}_{u}$.
@@ 3.1 Datasets
To provide guidelines for applying algorithms across diverse real-world applications, the selection of datasets in LLMNodeBed considers several key factors: (1) Multi-domain Diversity to reflect different contexts, (2) Varying Scales to examine algorithm scalability and the associated costs of leveraging LLMs, and (3) Diverse Levels of Homophily to understand its impact on performance. Therefore, LLMNodeBed comprises ten datasets spanning the academic, web link, social, and E-Commerce domains. These datasets vary significantly in scale, ranging from thousands of nodes to millions of edges, and exhibit differing levels of homophily. Such diversity in domain, scale, and homophily enables the assessment of algorithms across a wide range of contexts.
For datasets where raw text has been preprocessed into vector embeddings using bag-of-words or TF-IDF techniques, we utilize collected versions including Cora and Pubmed (He et al., 2023), Citeseer (Chen et al., 2024b), and WikiCS (Liu et al., 2024b). The remaining datasets already include text attributes in their official releases, including arXiv (Hu et al., 2020), Instagram and Reddit (Huang et al., 2024), and Books, Computer, and Photo (Yan et al., 2023). Detailed statistics and information for these datasets are provided in Table 8 and Table 9 in the Appendix.
@@ 3.2 Baselines
LLM-as-Encoder: We include ENGINE (Zhu et al., 2024b) and introduce GNNLLMEmb. ENGINE aggregates hidden embeddings from each LLM layer to create comprehensive node representations. In contrast, GNNLLMEmb initializes node embeddings using the LLM’s last hidden layer before feeding them into GNNs for classification.
LLM-as-Reasoner: We select TAPE (He et al., 2023), a representative LLM-as-Reasoner method. TAPE prompts the LLM to reason over nodes by generating predictions along with explanations, thereby enriching the node’s text attributes and enhancing classification performance.
Both Encoder and Reasoner methods require processing the entire dataset, either by encoding each node’s text or by performing reasoning over nodes, which introduces additional processing time before the actual model training begins.
@@
Besides LLM-based methods, LLMNodeBed also integrates classic algorithms, including MLPs, GNNs, and LMs. MLPs generate predicted label matrices from node embeddings, while GNNs combine shallow embeddings with graph structures for label prediction. LMs process a node’s text through hidden layers and use a classification head to produce label distributions. Further discussions of existing node classification algorithms are provided in Appendix A.
@@
Reasoner TAPE 84.04±${}_{\pm\text{0.24}}$ 71.87±${}_{\pm\text{0.35}}$ 78.61±${}_{\pm\text{1.23}}$ 81.94±${}_{\pm\text{0.16}}$ 66.07±${}_{\pm\text{0.10}}$ 62.43±${}_{\pm\text{0.47}}$ 84.92±${}_{\pm\text{0.26}}$ 86.46±${}_{\pm\text{0.12}}$ 89.52±${}_{\pm\text{0.04}}$ 78.43
@@
Reasoner TAPE 88.05±${}_{\pm\text{1.76}}$ 76.45±${}_{\pm\text{1.60}}$ 93.00±${}_{\pm\text{0.13}}$ 74.96±${}_{\pm\text{0.14}}$ 87.11±${}_{\pm\text{0.66}}$ 68.11±${}_{\pm\text{0.54}}$ 66.22±${}_{\pm\text{0.83}}$ 85.95±${}_{\pm\text{0.59}}$ 87.72±${}_{\pm\text{0.28}}$ 90.46±${}_{\pm\text{0.18}}$ 81.80
@@ 3.3 Learning Paradigms
  • Zero-shot Learning: No labeled data is provided for training. The model predicts labels solely based on node textual descriptions and the graph structure, assessing its ability to generalize to new, unseen data. For test samples, we follow existing literature (Zhu et al., 2024a) by selecting one smaller dataset from each domain and using 20% of its nodes as test samples.
@@ 4.1 Semi-supervised and Supervised Performance
Takeaway 1: Appropriately incorporating LLMs consistently improves the performance. According to the table, the best performance is often achieved by LLM-based methods compared to classic methods. It suggests that using LLM to exploit the textual information is useful.
@@
Takeaway 3: LLM-as-Reasoner methods are highly effective when labels heavily depend on text. TAPE achieves top or runner-up performance on academic and web link datasets like Cora and WikiCS, where structural information is less relevant to labels (Zhang et al., 2022b). However, TAPE struggles with social networks that require deeper structural understanding, such as predicting popular users (high-degree nodes) on Reddit.
Takeaway 4: LLM-as-Encoder methods balance computational cost and accuracy effectively. LLM-as-Encoder methods perform satisfactorily across all datasets. Further experiments in Section 5.1 reveal that LLM-as-Encoder methods are more effective than their LM counterparts when graphs are heterophilic. Regarding cost-effectiveness, LLM-as-Reasoner should generate long reasoning text, which is far more time-consuming than encoding texts in LLM-as-Encoder (see Appendix F). Therefore, LLM-as-Encoder methods strike a balance between computational efficiency and accuracy.
Takeaway 5: LLM-as-Predictor methods are more effective when labeled data is abundant. In supervised scenarios, LLM-as-Predictor methods enhance performance across most datasets. Especially, the LLaGA method achieves superior results among 5 of 10 datasets. Conversely, in semi-supervised settings, LLM-as-Predictor methods exhibit unstable performance, evidenced by low Macro-F1 scores, and imbalanced output distributions (detailed discussion in Appendix E.3). These findings indicate that LLM-as-Predictor methods are most effective when ample supervision is available, with LLaGA being an especially strong choice. Furthermore, within the predictor methods, LLM Instruction Tuning typically falls behind the other two methods and incurs substantial time costs (Table 17 and Table 18 in the Appendix). This shows that standalone LLMs are weak predictors and incorporating graph context is essential for achieving satisfactory performance.
4.2 Zero-shot Performance
Settings: For LLM Direct Inference, we utilize both closed-source and open-source LLMs, including GPT-4o (OpenAI, 2024), DeepSeek-Chat (Liu et al., 2024a), LLaMA3.1-8B (Dubey et al., 2024), and Mistral-7B. The prompt templates include Direct, CoT, ToT, and ReAct (Yao et al., 2023b). Additionally, we incorporate a node’s neighboring information into extended prompts, referred to as “w. Neighbor”, and have the LLMs first reason over neighbors to generate a summary that facilitates the subsequent classification task, referred to as “w. Summary”. Prompt templates are listed in Appendix B.2. Besides, we assess the transferability of GFMs by evaluating ZeroG (Li et al., 2024b), LLM Instruction Tuning, and LLaGA. GFMs are applied following the intra-domain manner: each model is pre-trained on a larger dataset within the same domain (e.g., arXiv from the academic domain) before being evaluated on the target dataset. Results for Accuracy and Macro-F1 are shown in Table 2 and Table 11 in the Appendix, where we have the following takeaway:
Table 2: Performance comparison under zero-shot setting with Accuracy (%) and Macro-F1 (%) reported.
@@ 4.2 Zero-shot Performance
Takeaway 6: GFMs can outperform open-source LLMs but still fall short of strong LLMs like GPT-4o. ZeroG outperforms LLaMA-8B in most cases, achieving up to a 6% average improvement in accuracy. However, it still falls short of GPT-4o and DeepSeek-Chat. Among GFMs, LLaGA performs poorly because it uses a projector to align the source graph’s tokens with LLM input tokens. This projector may be dataset-specific, leading to reduced performance on different datasets, as also observed in Zhu et al. (2024a). These findings highlight the need for further research to improve the generalization of GFMs to match the performance of more powerful LLMs.
Takeaway 7: LLM direct inference can be improved by appropriately incorporating structural information. Our results reveal that advanced prompt templates such as CoT, ToT, and ReAct, offer only minor performance improvements. Specifically, models like LLaMA exhibit limited instruction-following abilities, often producing unexpected and over-length outputs when processing complex prompts such as ReAct (Wu et al., 2025). This makes parsing classification results challenging and leads to suboptimal performance. The advanced prompts are generally designed for broad reasoning tasks and lack graph- or classification-specific knowledge, thereby limiting their benefits for the node classification task. In contrast, enriched prompts that incorporate structural information, i.e., “w. Neighbor” and “w. Summary”, demonstrate performance enhancements across LLMs. The performance gains are particularly evident on homophilic datasets such as Cora and Photo (3%-10%), where neighboring nodes are likely to share the same labels as the central node. High homophily means that information from neighboring nodes provides crucial clues about a central node’s label, thereby improving classification performance. Among these enriched prompts, “w. Summary” is especially effective as it not only provides structural context but also leverages the self-reflection abilities of LLMs to further utilize structural information.
Table 3: Comparison of LLM- and LM-as-Encoder with Accuracy (%) reported under semi-supervised setting. LLM-as-Encoder outperforms LMs in heterophilic graphs.

### v2

# When Do LLMs Help With Node Classification? A Comprehensive Analysis

@@ Abstract
Node classification is a fundamental task in graph analysis, with broad applications across various fields. Recent breakthroughs in Large Language Models (LLMs) have enabled LLM-based approaches for this task. Although many studies demonstrate the impressive performance of LLM-based methods, the lack of clear design guidelines may hinder their practical application. In this work, we aim to establish such guidelines through a fair and systematic comparison of these algorithms. As a first step, we developed LLMNodeBed, a comprehensive codebase and testbed for node classification using LLMs. It includes 10 homophilic datasets, 4 heterophilic datasets, 8 LLM-based algorithms, 8 classic baselines, and 3 learning paradigms. Subsequently, we conducted extensive experiments, training and evaluating over 2,700 models, to determine the key settings (e.g., learning paradigms and homophily) and components (e.g., model size and prompt) that affect performance. Our findings uncover 8 insights, e.g., (1) LLM-based methods can significantly outperform traditional methods in a semi-supervised setting, while the advantage is marginal in a supervised setting; (2) Graph Foundation Models can beat open-source LLMs but still fall short of strong LLMs like GPT-4o in a zero-shot setting. We hope that the release of LLMNodeBed, along with our insights, will facilitate reproducible research and inspire future studies in this field. Codes and datasets are released at https://llmnodebed.github.io/.
@@ 1 Introduction
Given its practical importance, node classification has been a long-standing research focus in both academia and industry. The earliest attempts to address this task adopted techniques such as Laplacian regularization (Belkin et al., 2006), graph embeddings (Yang et al., 2016), and label propagation (Zhu et al., 2003). Over the past decade, GNN-based methods have been developed and have quickly become prominent due to their superior performance, as demonstrated by works such as Kipf & Welling (2017), Veličković et al. (2018), and Hamilton et al. (2017). Additionally, the incorporation of encoded textual information has been shown to further complement GNNs’ node features, enhancing their effectiveness (Jin et al., 2023b; Zhao et al., 2023a).
@@
Inspired by the recent success of LLMs, there has been a surge of interest in leveraging LLMs for node classification (Li et al., 2023b). LLMs, pre-trained on extensive text corpora, possess context-aware knowledge and superior semantic comprehension, overcoming the limitations of the non-contextualized shallow embeddings used by traditional GNNs. Typically, supervised methods fall into three categories: Encoder, Explainer, and Predictor. In the Encoder paradigm, LLMs employ their vast parameters to encode nodes’ textual information, producing more expressive features that surpass shallow embeddings (Zhu et al., 2024). The Explainer approach utilizes LLMs’ generative capability to enhance node attributes and the task descriptions with a more detailed text (Chen et al., 2024b; He et al., 2024). This generated text augments the nodes’ original information, thereby enriching their attributes. Lastly, the Predictor role involves LLMs integrating graph context through graph encoders, enabling direct text-based predictions (Chen et al., 2024a; Tang et al., 2023; Chai et al., 2023; Huang et al., 2024). For zero-shot learning with LLMs, methods can be categorized into two types: Direct Inference and Graph Foundation Models (GFMs). Direct Inference involves guiding LLMs to directly perform classification tasks via crafted prompts (Huang et al., 2023). In contrast, GFMs entail pre-training on extensive graph corpora before applying the model to target graphs, thereby equipping the model with specialized graph intelligence (Li et al., 2024b). An illustration of these methods is shown in Figure 2.
@@
Despite tremendous efforts and promising results, the design principles for LLM-based node classification algorithms remain elusive. Given the significant training and inference costs associated with LLMs, practitioners may opt to deploy these algorithms only when they provide substantial performance enhancements compared to costs. This study, therefore, seeks to identify (1) the most suitable settings for each algorithm category, and (2) the scenarios where LLMs surpass traditional LMs such as BERT. While recent work like GLBench (Li et al., 2024c) has evaluated various methods using consistent data splits in semi-supervised and zero-shot settings, differences in backbone architectures and implementation codebases still hinder fair comparisons and rigorous conclusions. To address these limitations, we introduce a new benchmark that further standardizes backbones and codebases. Additionally, we extend GLBench by incorporating 3 E-Commerce datasets, and 4 heterophilic datasets, while also expanding the evaluation settings. Specifically, we assess the impact of supervision signals (e.g., supervised, semi-supervised), different language model backbones (e.g., RoBERTa, Mistral, LLaMA, Qwen, GPT-4o), and various prompt types (e.g., CoT, ToT, ReAct). These enhancements enable a more detailed and reliable analysis of LLM-based node classification methods. In summary, our contributions to the field of LLMs for graph analysis are as follows:
@@
  • A Testbed: We release LLMNodeBed, a PyG-based testbed designed to facilitate reproducible and rigorous research in LLM-based node classification algorithms. The initial release includes 14 datasets, 8 LLM-based algorithms, and 3 learning configurations. LLMNodeBed allows for easy addition of new algorithms or datasets, and a single command to run all experiments, and to automatically generate all tables included in this work.
@@
  • Comprehensive Experiments: By training and evaluating over 2,700 models, we analyzed how the learning paradigm, homophily, language model type and size, and prompt design impact the performance of each algorithm category.
@@
Table 1: Statistics of supported datasets in LLMNodeBed.
Academic Web Link Social E-Commerce Heterophilic
Statistics Cora Citeseer Pubmed arXiv WikiCS Instagram Reddit Books Photo Computer Cornell Texas Wisconsin Washington
# Classes 7 6 3 40 10 2 2 12 12 10 5 5 5 5
# Nodes 2,708 3,186 19,717 169,343 11,701 11,339 33,434 41,551 48,362 87,229 191 187 265 229
# Edges 5,429 4,277 44,338 1,166,243 215,863 144,010 198,448 358,574 500,928 721,081 292 310 510 394
Avg. # Token 183.4 210.0 446.5 239.8 629.9 56.2 197.3 337.0 201.5 123.1 594.6 453.2 639.1 469.0
Homophily (%) 82.52 72.93 79.24 63.53 68.67 63.35 55.52 78.05 78.50 85.28 11.55 6.69 16.27 17.07
@@ 2 Preliminaries on Node Classification
To leverage the language abilities of LLMs, we study the node classification task within the context of text-attributed graphs (TAGs) (Ma & Tang, 2021). TAGs are represented as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{S})$, where $\mathcal{V}$ represents the set of nodes, $\mathcal{E}$ the set of edges, and $\mathcal{S}$ the collection of textual descriptions associated with each node $v\in\mathcal{V}$. Some of the nodes are associated with labels, represented as $\mathcal{V}_{l}\subset\mathcal{V}$. The remaining nodes do not have labels, and are denoted as $\mathcal{V}_{u}$. The goal of node classification is to train a neural network based on the graph $\mathcal{G}$ and the labels of $\mathcal{V}_{l}$, which can predict the labels of unlabeled nodes in $\mathcal{V}_{u}$.
@@ 3.1 Datasets
To provide guidelines for applying algorithms across diverse real-world applications, the selection of datasets in LLMNodeBed considers several key factors: (1) Multi-domain Diversity to reflect different contexts, (2) Varying Scales to examine algorithm scalability and the associated costs of leveraging LLMs, and (3) Diverse Levels of Homophily to understand its impact on performance. Therefore, LLMNodeBed comprises $14$ datasets spanning various domains. These datasets vary significantly in scale, ranging from thousands of nodes to millions of edges, and exhibit differing levels of homophily. Such diversity in domain, scale, and homophily enables the assessment of algorithms across a wide range of contexts.
For datasets where raw text has been preprocessed into vector embeddings using bag-of-words or TF-IDF techniques, we utilize collected versions including Cora and Pubmed (He et al., 2024), Citeseer (Chen et al., 2024b), and WikiCS (Liu et al., 2024). The remaining datasets already include text attributes in their official releases, including arXiv (Hu et al., 2020), Instagram and Reddit (Huang et al., 2024), Books, Computer, and Photo (Yan et al., 2023), Cornell, Texas, Wisconsin, and Washington (Wang et al., 2025). Detailed statistics and information for these datasets are provided in Table 1 and Table 10 in the Appendix.
@@ 3.2 Baselines
LLM-as-Encoder: We include ENGINE (Zhu et al., 2024) and introduce GNNLLMEmb. ENGINE aggregates hidden embeddings from each LLM layer to create comprehensive node representations. In contrast, GNNLLMEmb initializes node embeddings using the LLM’s last hidden layer before feeding them into GNNs for classification.
LLM-as-Explainer: We select TAPE (He et al., 2024), a representative LLM-as-Explainer method. TAPE prompts the LLM to think over nodes by generating predictions along with explanations, thereby enriching the node’s text attributes and enhancing classification performance.
Both Encoder and Explainer methods require processing the entire dataset, either by encoding each node’s text or generating explanations for nodes, which introduces additional processing time before the actual model training begins.
@@
Besides LLM-based methods, LLMNodeBed also integrates classic algorithms, including MLPs, GNNs, and LMs. MLPs generate predicted label matrices from node embeddings, while GNNs combine shallow embeddings with graph structures for label prediction. LMs process a node’s text through hidden layers and use a classification head to produce label distributions. Furthermore, we include GLEM (Zhao et al., 2023a), a hybrid LM+GNN method that combines graph and text modalities. Further discussions of existing algorithms are provided in Appendix A.
@@
GLEM 81.30±${}_{\pm\text{0.88}}$ 68.80±${}_{\pm\text{2.46}}$ 81.70±${}_{\pm\text{1.07}}$ 76.43±${}_{\pm\text{0.55}}$ 60.25±${}_{\pm\text{3.66}}$ 55.13±${}_{\pm\text{1.41}}$ 83.28±${}_{\pm\text{0.39}}$ 76.93±${}_{\pm\text{0.49}}$ 80.46±${}_{\pm\text{1.45}}$ 73.81
@@
Explainer TAPE 84.04±${}_{\pm\text{0.24}}$ 71.87±${}_{\pm\text{0.35}}$ 78.61±${}_{\pm\text{1.23}}$ 81.94±${}_{\pm\text{0.16}}$ 66.07±${}_{\pm\text{0.10}}$ 62.43±${}_{\pm\text{0.47}}$ 84.92±${}_{\pm\text{0.26}}$ 86.46±${}_{\pm\text{0.12}}$ 89.52±${}_{\pm\text{0.04}}$ 78.43
@@
GLEM 86.81±${}_{\pm\text{1.19}}$ 73.24±${}_{\pm\text{1.55}}$ 93.98±${}_{\pm\text{0.32}}$ 73.55±${}_{\pm\text{0.22}}$ 79.81±${}_{\pm\text{0.45}}$ 67.39±${}_{\pm\text{1.73}}$ 53.11±${}_{\pm\text{2.96}}$ 83.98±${}_{\pm\text{0.97}}$ 78.16±${}_{\pm\text{0.45}}$ 81.63±${}_{\pm\text{0.46}}$ 77.17
@@
Explainer TAPE 88.05±${}_{\pm\text{1.76}}$ 76.45±${}_{\pm\text{1.60}}$ 93.00±${}_{\pm\text{0.13}}$ 74.96±${}_{\pm\text{0.14}}$ 87.11±${}_{\pm\text{0.66}}$ 68.11±${}_{\pm\text{0.54}}$ 66.22±${}_{\pm\text{0.83}}$ 85.95±${}_{\pm\text{0.59}}$ 87.72±${}_{\pm\text{0.28}}$ 90.46±${}_{\pm\text{0.18}}$ 81.80
@@ 3.3 Learning Paradigms
  • Zero-shot Learning: No labeled data is provided for training. The model predicts labels solely based on node textual descriptions and the graph structure, assessing its ability to generalize to new, unseen data. For test samples, we follow existing literature (Zhu et al., 2025) by selecting one smaller dataset from each domain and using 20% of its nodes as test samples.
@@ 4.1 Semi-supervised and Supervised Performance
Takeaway 1: Appropriately incorporating LLMs consistently improves the performance. According to the table, the best performance is often achieved by LLM-based methods compared to classic methods. It suggests that using LLM to exploit the textual information is useful. This conclusion extends to large-scale graphs as well, e.g., ogbn-product dataset with millions of nodes. Detailed results and analysis for this dataset are provided in Appendix D.
@@
Takeaway 3: LLM-as-Explainer methods are highly effective when labels heavily depend on text. TAPE achieves top or runner-up performance on academic and web link datasets like Cora and WikiCS, where structural information is less relevant to labels (Zhang et al., 2022b). However, TAPE struggles with social networks that require deeper structural understanding, such as predicting popular users (high-degree nodes) on Reddit.
Takeaway 4: LLM-as-Encoder methods balance computational cost and accuracy effectively. LLM-as-Encoder methods perform satisfactorily across all datasets. Further experiments in Section 5.1 reveal that LLM-as-Encoder methods are more effective than their LM counterparts when graphs are less informative. Regarding cost-effectiveness, LLM-as-Explainer should generate long explanatory text, which is far more time-consuming than encoding texts in LLM-as-Encoder (see Appendix G). Therefore, LLM-as-Encoder methods strike a balance between computational efficiency and accuracy.
Takeaway 5: LLM-as-Predictor methods are more effective when labeled data is abundant. In supervised scenarios, LLM-as-Predictor methods enhance performance across most datasets. Especially, the LLaGA method achieves superior results among 5 of 10 datasets. Conversely, in semi-supervised settings, LLM-as-Predictor methods exhibit unstable performance, evidenced by low Macro-F1 scores, and imbalanced output distributions (detailed discussion in Appendix F.3). These findings indicate that LLM-as-Predictor methods are most effective when ample supervision is available, with LLaGA being an especially strong choice. Furthermore, within the predictor methods, LLM Instruction Tuning typically falls behind the other two methods and incurs substantial time costs (Table 19 in the Appendix). This shows that standalone LLMs are weak predictors and incorporating graph context is essential for achieving satisfactory performance.
[…]
@@
4.2 Zero-shot Performance
Settings: For LLM Direct Inference, we utilize both closed-source and open-source LLMs, including GPT-4o (OpenAI, 2024), DeepSeek-V3 (DeepSeek-AI, 2024), LLaMA3.1-8B (Dubey et al., 2024), and Mistral-7B. The prompt templates include Direct, CoT, ToT, and ReAct (Yao et al., 2023b). Additionally, we incorporate a node’s neighboring information into extended prompts, referred to as “w. Neighbor”, and have the LLMs first reason over neighbors to generate a summary that facilitates the subsequent classification task, referred to as “w. Summary”. Prompt templates are listed in Appendix B.2. Besides, we assess the transferability of GFMs by evaluating ZeroG (Li et al., 2024b), LLM Instruction Tuning, and LLaGA. GFMs are applied following the intra-domain manner: each model is pre-trained on a larger dataset within the same domain (e.g., arXiv from the academic domain) before being evaluated on the target dataset. Results for Accuracy and Macro-F1 are shown in Table 3 and Table 13 in the Appendix, where we have the following takeaways:
Takeaway 6: GFMs can outperform open-source LLMs but still fall short of strong LLMs like GPT-4o. ZeroG outperforms LLaMA-8B in most cases, achieving up to a 6% average improvement in accuracy. However, it still falls short of GPT-4o and DeepSeek-V3. Among GFMs, LLaGA performs poorly because it uses a projector to align the source graph’s tokens with LLM input tokens. This projector may be dataset-specific, leading to reduced performance on different datasets, as also observed in Zhu et al. (2025). These findings highlight the need for further research to improve the generalization of GFMs to match the performance of more powerful LLMs.
Takeaway 7: LLM direct inference can be improved by appropriately incorporating structural information. Our results reveal that advanced prompt templates such as CoT, ToT, and ReAct, offer only minor performance improvements. Specifically, models like LLaMA exhibit limited instruction-following abilities, often producing unexpected and over-length outputs when encountering complex prompts such as ReAct. This makes parsing classification results challenging and leads to suboptimal performance. The advanced prompts are generally designed for broad reasoning tasks and lack graph- or classification-specific knowledge, thereby limiting their benefits for the node classification task. In contrast, enriched prompts that incorporate structural information, i.e., “w. Neighbor” and “w. Summary”, demonstrate performance enhancements across LLMs. The performance gains are particularly evident on homophilic datasets such as Cora and Photo (3%-10%), where neighboring nodes are likely to share the same labels as the central node. High homophily means that information from neighboring nodes provides crucial clues about a central node’s label, thereby improving classification performance. Among these enriched prompts, “w. Summary” is especially effective as it not only provides structural context but also leverages the self-reflection abilities of LLMs to further utilize structural information.
[…]